OpenStudy (anonymous):

Identify the slope and the y-intercept for the line equation given: −5x + y = −3.

OpenStudy (anonymous):

what you need to do is put your equation in y = mx + b form where m is your slope and b is your y intercept. Can you do this ?

OpenStudy (anonymous):

Not exactly.

OpenStudy (anonymous):

-5x + y = - 3 --- add 5x to both sides -5x + 5x + y = -3 + 5x -- combine like terms y = 5x - 3 this is now in y = mx + b form. Your slope is in the m position...so what is your slope ? Your y intercept is in the b position...so what is your y intercept ?

OpenStudy (anonymous):

any thoughts ?

OpenStudy (anonymous):

I have no idea how to do this.

OpenStudy (anonymous):

 as you can see, your slope is 5 and your y intercept is -3